In a mixed class, there are 10 girls.
If 2 pupils are selected at random, the probability that they are both girls is 0.15
How many boys are in the class?

Answer by stanbon(75887)

In a mixed class, there are 10 girls.
If 2 pupils are selected at random, the probability that they are both girls is 0.15
How many boys are in the class?
-------------------------
Let # of boys be "x".
-----
0.15 = 10C2/(10+x)C2
---
0.15 = (45)/[(10+x)(9+x)]/2
--------------------------
0.15 = 90/[(10+x)(9+x)]
-------
1/600 = 1/[(10+x)(9+x)]
-------
(10+x)(9+x) = 600
-------
x^2+19x+90 = 600
----
x^2 + 19x - 510 = 0
---
Factor:
(x-15)(x+34) = 0
----
Positive solution:
x = 15 (# of boys in the class)
=====================================
Cheers,
Stan H.